OpenStudy (anonymous):

Simplify: 16 < 8(x - 2) A. x > 1/2 B. 4 < x C. x > 2 D. x < -4 @sangya21

OpenStudy (anonymous):

16<8x - 16 add 16 to both sides = 16+16 <8x and then divide by 8 = 32/8 <x 4<x

OpenStudy (anonymous):

16 <8(x - 2) 16 < 8x -16 16 + 16 <8x 32 < 8x 32/8 <x 4<X

OpenStudy (radar):

16<8(x-2) divide both sides by 8 2< x-2 add 2 to both sides 4<x Just another way to proceed.